OpenStudy (anonymous):

Let f(x) = 27x5 - 33x4 - 21x3 and g(x) = 3x2. Find f of x over g of x 9x2 - 11x - 7 9x2 + 11x + 7 9x3 - 11x2 - 7x 9x3 + 11x2 + 7x

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

here is a common factor of 3x^2 in each term leaving 9x^3 - 11x^2 - 7x

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

so c @burn23

OpenStudy (anonymous):

yes C

OpenStudy (anonymous):

can you help we couple more :)

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

Divide 3x2 - 11x - 4 by x - 4. x + 1 x - 23 3x + 1 3x - 23

OpenStudy (anonymous):

3x2–11x–4/ x–4=(x−4)(3x+1)/x−4=3x+1

OpenStudy (anonymous):

Let f(x) = 8x3 - 28x + 61 and g(x) = 2x + 5. Find f of x over g of x. 4x2 + 10x + 34 + 116 over quantity of 2 x plus 5 4x2 + 10x + 34 + quantity of 2 x plus 5 over 116 4x2 - 10x + 11+ quantity of 2 x plus 5 over 6 4x2 - 10x + 11+ 6 over quantity of 2 x plus 5

OpenStudy (anonymous):

@burn23

OpenStudy (anonymous):

4x^2-10x+11+6/2x+5

OpenStudy (anonymous):

Divide 36x3 - 96x2 + 48x by 12x. 3x2 - 8x + 4 3x2 + 8x - 4 3x3 - 8x + 4 3x3 + 8x - 4

OpenStudy (anonymous):

@burn23